A researcher was interested in whether there was differences in the effect on performance on a stats exam among students who attended

1. Office hours regularly,

2. Both office hours and review sessions.

3. Just review sessions, and

4. Neither office hours nor review sessions.

The researcher randomly assigned students to one of the four groups. At the end of the first exam, the researcher gave students in all four groups a stats exam and obtained the following results:

Office Office and Review Review None
12 12 8 6
11 10 7 9
14 15 10 11
11 14 9 10
12 14 6 9
x with a bar O=12,ssO=6 x with bar O&R=13,SSO&R=16 x with a bar R=s, SSR=10 x with a bar N=9, SSN=14

The researcher used an a-level of .05 to test her hypothesis. I want to see four steps, step 1, formulate hypotheses; step 2, select criteria (α, dfB, dfW, critical F); step 3, calculation; step 4, conclusion (we reject or fail to reject null hypothesis and infer that … (F(dfB, dfW)= , p )).
Answer #1

(1) Ho: All four groups have the same mean (μ1 = μ2 = μ3 = μ4)

Ha: At least two of the four groups have significantly different means

(2) α = 0.05, dfB = 4 - 1 = 3, dfT = 20 - 1 = 19, dfW = 19 - 3 = 16

(3 and 4)

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]